Golf ball exploration!
I am doing this project for class, but I am having a little difficulty can you help me solve the problem?
Directions: Explore the relationship between the height of a golf ball before dropping it and its height after one bounce.
*THIS IS WHAT IT WANTS ME TO DO: A golf ball should be dropped from an initial height of 12 inches. (NOTE: I am dropping this BALL on a smooth concrete pavement) Estimate the bounce height of the ball. Then do it again only this time drop it from 24 inches, 36 inches, 48 inches, 60 and 72 inches. Then compute the average bounce heights for each of the drops. Then turn it into an equation. How do I do this?
1
Expert's answer
2011-11-29T11:15:00-0500
When I have droped it from 12 inches I received 8 height after one bounce
When I have droped it from 24 inches I received 16 height after one bounce
When I have droped it from 36 inches I received 24 height after one bounce
When I have droped it from 48 inches I received 32 height after one bounce
When I have droped it from 60 inches I received 40 height after one bounce
When I have droped it from 72 inches I received 48 height after one bounce
So the formula should be h=12/8x=3/2x or x=2/3h,
Where h- height of a golf ball before dropping
And x- height after one bounce